OpenStudy (anonymous):

Will Medal And Fan IF Given Correct Answer. Find the slope and the y-intercept of the equation 4x + 5y = -9.

Directrix (directrix):

The task is to write 4x + 5y = -9 in y = mx+b form

Directrix (directrix):

4x + 5y = -9 5y = -4x -9

OpenStudy (anonymous):

No dude i ave to find the slope and y-intercept

Directrix (directrix):

5y = -4x -9 Divide each term by 5. y = (-4/5) x - 9/5

Directrix (directrix):

I understand that.

OpenStudy (alexandervonhumboldt2):

-4/5 slope -9/5 y-inter

Directrix (directrix):

y = (-4/5) x - 9/5 y = m x + b m = -4/5, y-intercept = -9/5. The y intercept may be shown as an ordered pair (0, -9/5) @ALG3BRUHH69

Directrix (directrix):

@AlexandervonHumboldt2 Posting answers with no work is a violation of the OpenStudy Code of Conduct.

Directrix (directrix):

@ALG3BRUHH69

OpenStudy (anonymous):

I best response you, that is giving you a medal isnt it ?

Directrix (directrix):

Yes it is. Thank you.